# ENTROPY-REGULARIZED GRADIENT ESTIMATORS FOR APPROXIMATE BAYESIAN INFERENCE

### v2

@@ Abstract
Effective uncertainty quantification is important for training modern predictive models with limited data, enhancing both accuracy and robustness. While Bayesian methods are effective for this purpose, they can be challenging to scale. When employing approximate Bayesian inference, ensuring the quality of samples from the posterior distribution in a computationally efficient manner is essential. This paper addresses the estimation of the Bayesian posterior to generate diverse samples by approximating the gradient flow of the Kullback-Leibler (KL) divergence and the cross entropy of the target approximation under the metric induced by the Stein Operator. It presents empirical evaluations on classification tasks to assess the method’s performance and discuss its effectiveness for Model-Based Reinforcement Learning that uses uncertainty-aware network dynamics models.
@@ 1 Introduction
Deep Learning models often result in overconfident and miscalibrated predictions without uncertainty quantification. It becomes more important when limited data is available, the quality of data available is poor, or one needs to consider making out-of-distribution predictions. These cases find their use in real-world situations that require safe decision-making. Measuring uncertainty in predictions plays a crucial role in Reinforcement Learning (RL). Many works incorporate uncertainty to guide exploration, directing it toward regions of high uncertainty (e.g., Lee et al. [2021], Liu et al. [2021], Ciosek et al. [2019], Osband et al. [2016]). Chua et al. [2018], Fu et al. [2022], and Janner et al. [2019] utilize ensembles of models to capture and represent uncertainty in learned environment dynamics, employing these models for planning purposes. In other cases, deep ensembles effectively mitigate overestimation bias in Q-value functions and enhance stability (Chen et al. [2021]).
@@
Methods using ensembles have become popular for uncertainty modeling in deep neural networks. Ensembles combine several individual predictors. An ensemble’s predictive performance depends on individual performance, and the diversity of individual members plays an important role. Recent work, Abe et al. [2023], Abe et al. [2022a], points out how the diversity of the members gets reduced as the networks get larger. The effectiveness of uncertainty quantification using an ensemble of deep neural networks has been shown to reduce as the predictive ability of the ensemble members increases. This can be attributed to having large enough in-distribution data, big ensemble size, or over-parameterized networks. Having training objectives that train an ensemble to make diverse predictions may seem like a possible solution, but this can harm the predictive performance (Abe et al. [2023]). This has driven recent research to understand the cases under which ensembles are truly effective. A good single model can outweigh the benefits of using an ensemble for making out-of-distribution predictions ( Abe et al. [2022b]). However, diversity is still important for predictions in a limited data regime and can not be easily mitigated by increasing the model size without overfitting.
@@
Work in recent years has explored methods for Bayesian inference to obtain predictive ensembles and has demonstrated efficacy in uncertainty estimation (Neal [2012], Mackay [1992]). In Bayesian inference, one usually cares about the distribution of model parameters. Once this distribution is estimated, it can be used to make predictions by marginalizing. The way to approach this in a tractable manner is referred to as Bayesian Model Averaging. Such a class of methods has been shown to improve generalization, enhance uncertainty quantification, and improve model reliability in handling out-of-distribution data.
@@
Traditional approaches to Bayesian inference depend on obtaining samples from Markov Chain Monte Carlo (MCMC) chains. These become computationally inefficient for bigger networks. Some approximate methods, such as stochastic gradient MCMC methods, manage to generate samples of the posterior distribution with the help of gradient information and use mini-batches of data. In comparison to these, Explicit variational methods are significantly better at inference time. This is because they directly optimize a tractable approximation of the true posterior, as opposed to depending on slower sampling methods like SG-MCMC. Implicit Variational methods, also known as Predictive Modeling, focus on directly estimating approximate posterior samples and have shown strong empirical performance. Posterior samples obtained using implicit methods constitute an ensemble of parameterized networks.
@@
This paper considers a type of Implicit method: Particle optimization methods for approximate Bayesian Inference, which approximate the target distribution with an ensemble of networks acting as a set of particles. One common issue with these methods is that they often experience mode collapse. This occurs when the different members of the ensemble make similar predictions. Having functional diversity in posterior samples makes it easy to avoid averaging over similar functions, which leads to redundancy in the model average (D’Angelo and Fortuin [2021b], Wu et al. [2017], Pang et al. [2019]). This phenomenon of mode collapse is more evident when networks are over-parameterized (D’Angelo and Fortuin [2021b], Kirsch [2024]). It has been described as the problem of non-identifiability (Roeder et al. [2021]).
@@
This work addresses the mode collapse issue and proposes a solution to ensure functional diversity using gradient information. It examines how functional diversity in particle optimization methods can be effectively controlled with kernel density estimation to approximate the target density. The paper introduces a simple method for training ensemble members for improved mode exploration and presents empirical results to validate the approach.
@@ 1.1 Contribution
Empirically study the implications of the method and test its efficacy in improving planning using MPC for Model-Based RL.
2 Preliminaries
@@ 2.1 Bayesian Neural Networks (BNNs)
In a supervised learning setting, one has where denoting the training inputs and denote the corresponding outputs, with and , respectively. Let denote a mapping function parameterized by a neural network, and one can define a likelihood function using this .
@@
In Bayesian Neural Networks, the aim is to estimate the posterior distribution that captures all possible explanations for the given training data, defined as follows.
represents the prior distribution over the weight parameters . This helps capture a type of uncertainty called epistemic uncertainty. Epistemic uncertainty modeling especially shines when data is limited or additional data collection is difficult. Although many works use epistemic uncertainty methods when distributional shifts are present, this could be potentially misleading. Ensembles are essential for improving predictive performance, uncertainty quantification, and robustness over a single model. However, their ability to quantify uncertainty and ensure robustness on out-of-distribution data correlates with their in-distribution performance. Work such as Malinin and Gales [2018] describes a non-parametric way to quantify uncertainty for such distribution shifts.
Most traditional ways to use neural network methods use a point estimate called a Maximum A Posteriori (MAP) estimate, which can be useful when sampling from a Bayesian Neural Network is challenging or when no model uncertainty needs to be captured.
@@
However, this integral is intractable for neural networks. It is typically approximated using Bayesian Model Averaging (BMA). Sampling-based methods for Bayesian Inference facilitate the acquisition of representative samples from such a posterior distribution to estimate this quantity. This becomes the reason why prediction accuracy in most approximate methods relies heavily on this Bayesian Model Averaging estimate.
2.2 Particle-optimization-based Variational Inference
This section describes how Implicit methods estimate the posterior as a stationary distribution by minimizing an energy functional. This can be modeled as a partial differential equation of the form :
where is the approximate posterior at time , and is the gradient flow. This gradient flow depends on the choice of metric and energy functional. Samples obtained by simulating such a gradient flow constitute samples from this stationary distribution. In practice, numerical methods can be used for simulation. Particle-optimization based Variational Inference (POVI) methods approximate with a set of particles , i.e.,
and simulate the gradient flow with a discretized version of the ODE
@@ 2.2 Particle-optimization-based Variational Inference
In this way, implicit variational methods allow for the generation of samples without having to do probability evaluation for approximating the true posterior. Initial particle distribution plays a key role in the diversity of these samples. As the networks get larger, they become less and less sensitive to this initial distribution during training and converge to a single mode. This contrasts with Explicit variational methods, which approximate the posterior distribution by minimizing the KL divergence between a simpler, tractable distribution and the true posterior. Thus, the use of limited tractable distributions can hinder the accuracy of these methods.
2.3 Gradient Flows
Different methods use different gradient flows. This section describes some of the most common choices of metric spaces and functionals for the gradient flow in POVI.
The -Wasserstein distance for between and is defines a metric on as
It can be understood as an optimal transport problem, where one wants to transform elements in the domain of to with a minimum cost.
Considering the optimization problem of a functional , the evolution in time of the measure under the Wasserstein gradient flow is described by
as described in Jordan et al. [1998], Liu and Liu [2017] in detail where
@@ 2.3 Gradient Flows
Some examples of Particle-Optimization Sampling methods that use Wasserstein gradient flow are D’Angelo and Fortuin [2021b], which uses kernel density to approximate the target distribution. Chen et al. [2018] presents a novel particle-optimization framework grounded in Wasserstein gradient flows to unify two prominent scalable Bayesian sampling methods: stochastic gradient Markov chain Monte Carlo (SG-MCMC) and Stein variational gradient descent (SVGD).
D’Angelo and Fortuin [2021b] shows how SVGD can be reinterpreted as a gradient flow in the Wasserstein space under a particular Stein geometry induced on the manifold by the kernel and where is the metric. If is a linear operator defined as , the ODE describing the SVGD update becomes :
3 Deep Ensembles by ERGD
Empirical evidence and theoretical analysis Zhuo et al. [2018] have shown that particles in methods such as SVGD exhibit mode collapse. They converge to a limited number of local modes. Many works have proposed solutions to optimize for mode coverage by considering the impact of the interaction between the particles. This phenomenon is strongly related to the particles’ initial distribution and becomes worse as the dimensionality increases. This section introduces a method to enhance mode coverage by mitigating the ensemble’s dependence on the initial distribution while accounting for particle interactions within the ensemble.
The proposed method is called Entropy-Regularized Gradient Estimators. It uses an empirical measure of particles to approximate the target posterior. These set of particles are taken from an initial distribution and are evolved using the following update :
As in most similar works, a symmetric kernel models the interaction between the particles to train an ensemble. ERGD introduces a regularization constant that controls the effect of this interaction on functional diversity. The key idea behind this is that using kernel density as an approximation to the target density, the constant regularizes cross-entropy between the kernel density and this approximation. Intuitively, the evolution of the particles is influenced more by their interactions with each other.
@@ 3 Deep Ensembles by ERGD
The convergence guarantees of the method determine the choice for this regularization constant. As , is a solution of the corresponding Louiville Equation (more details in Appendix A). A simple choice for satisfying this is to use a linear schedule between positive value bigger than 1 to 1. This provides a method to guide particle evolution, placing much greater emphasis on the interactions between the particles in the initial distribution.
Using a symmetric kernel as an approximation to target probability density changes the gradient flow of KL divergence between two distributions under the metric induced by the Stein Operator to the gradient flow of KL divergence under the Wasserstein Metric (D’Angelo and Fortuin [2021b]). The following update gives the particle evolution method described in D’Angelo and Fortuin [2021b] :
It can be understood as having the ensemble members grow without attraction, leading to a high probability density. In this case, the authors used as 1 / . Such a can be bad in practice (D’Angelo and Fortuin [2021b], D’Angelo and Fortuin [2021b]). Kirsch [2024] developed gradient estimators by inverting Stein’s Identity. The work demonstrates that their method of selecting outperforms the gradient estimator proposed in D’Angelo and Fortuin [2021b], as their estimator accounts for the effects of all interactions on a particle, including those not involving the particle itself. ERGD can direct particle evolution based on an initial distribution, but it’s performance depends on this initial distribution. In order to find a distribution that optimizes for mode coverage, ERGD is extended to -ERGD as follows :
-ERGD update makes a tunable parameter in the algorithm. Experiments comparing different updates use for -ERGD. Although this update is similar to Eq(6), it does not guarantee convergence.
@@ 4.2 Classification Tasks
Different methods are compared for an ensemble size of 10 and 20 for the classification of CIFAR-10 and FMNIST, respectively. Classification on CIFAR-10 is done using ResNet 20 and Fashion-MNIST uses a neural network with 3 hidden layers and 128 hidden units. ERGD and -ERGD are compared with SGD, kde-WGD using an update rule given by [6], sse-WGD as described in Kirsch [2024] and SVGD. ERGD uses a linear annealing schedule. For -ERGD, experiments use a constant value of beta in order to compare with kde-WGD and sse-WGD.
ERGD performs better than SVGD and is comparative to kde-WGD and sse-WGD for both the datasets. Deep Ensembles perform comparative to kde-WGD on CIFAR-10 and have comparative performance to -ERGD on FMINST.
@@
4.3 Multi-Step Predictions for Model-Based RL
@@ 5 Related Work
In the context of POVI, making diverse predictions in weight-space does not mitigate the issue of non-identifiability in large networks, and function-space methods as described in D’Angelo and Fortuin [2021b] and Sun et al. [2019] perform better.
5.1 Conclusion

### v3

@@ Abstract
Effective uncertainty quantification is important for training predictive models with limited data to improve accuracy and robustness. While Bayesian methods are effective for this purpose, they can be challenging to scale. Ensuring the quality of samples from the posterior distribution in a computationally efficient manner becomes essential when employing approximate Bayesian inference. This paper addresses the estimation of the Bayesian posterior to generate diverse samples by approximating the gradient flow of the Kullback-Leibler (KL) divergence and the cross entropy of the target approximation under the metric induced by the Stein Operator. It presents empirical evaluations on classification tasks to assess the method’s performance and discuss its effectiveness for Model-Based Reinforcement Learning that uses uncertainty-aware network dynamics models.
@@ 1 Introduction
Deep Learning models often result in overconfident and miscalibrated predictions without uncertainty quantification. This is more pronounced in cases when limited data is available, the quality of data available is poor, or one needs to consider making out-of-distribution predictions. Moreover, such cases find their use in real-world situations that require safe decision-making. Measuring uncertainty in predictions plays a crucial role in Reinforcement Learning (RL). Many works incorporate uncertainty to guide exploration, directing it toward regions of high uncertainty (e.g., Lee et al. [2021], Liu et al. [2021], Ciosek et al. [2019], Osband et al. [2016]). Chua et al. [2018], Fu et al. [2022], and Janner et al. [2019] utilize ensembles of models to capture and represent uncertainty in learned environment dynamics, employing these models for planning purposes. In many cases, deep ensembles effectively mitigate overestimation bias and stabilize Q-value functions that quantify the value of an action in a given state (Chen et al. [2021]).
@@
Methods using ensembles have become popular for uncertainty modeling in deep neural networks. Ensembles combine several individual predictors. An ensemble’s predictive performance depends on individual performance and the diversity among the individual members. Recent work, Abe et al. [2023], Abe et al. [2022a], points out how the diversity of the members reduces as the networks get larger. In general, the effectiveness of uncertainty quantification using an ensemble of deep neural networks reduces as the predictive ability of the ensemble members increases. This can be attributed to having large enough in-distribution data, big ensemble size, or over-parameterized networks. Having training objectives that train an ensemble to make diverse predictions may seem like a possible solution, but this can harm the predictive performance (Abe et al. [2023]). This led to research to understand the cases under which ensembles are truly effective. A good single model can outweigh the benefits of using an ensemble for making out-of-distribution predictions ( Abe et al. [2022b]). However, diversity is still important for predictions in a limited data regime and can not be easily mitigated by increasing the model size without overfitting.
@@
Recent work builds on the contributions of Neal [2012] and Mackay [1992], extending Bayesian inference methods to develop predictive ensembles and improve uncertainty estimation. In Bayesian inference, one usually cares about the distribution of model parameters. Once this distribution is estimated, it can be used to make predictions by marginalizing. Bayesian Model Averaging is a way to do this in a tractable manner, with samples of this distribution constituting an ensemble. Such a class of methods has been shown to improve generalization, enhance uncertainty quantification, and improve model reliability in handling out-of-distribution data.
@@
Traditional approaches to Bayesian inference depend on obtaining samples from Markov Chain Monte Carlo (MCMC) chains. And, these become computationally inefficient for bigger networks. Some approximate methods, such as stochastic gradient MCMC methods, manage to generate samples of the posterior distribution with the help of gradient information and use mini-batches of data. In comparison to these slow sampling methods, methods that directly optimize a tractable approximation of the true posterior tend to do better at inference time. These are often referred to as Explicit Variational Methods. Implicit Variational methods, on the other hand, focus on directly estimating approximate posterior samples and have shown strong empirical performance (li2017gradient).
@@
This paper considers a type of Implicit method: Particle optimization methods for approximate Bayesian Inference, which approximate the target distribution with an ensemble of networks acting as a set of particles and improve diversity among them. More specifically, this paper deals with the issue of mode collapse that occurs when the different members of the ensemble make similar predictions. It introduces an approach to have functional diversity in posterior samples to avoid averaging over similar functions, which leads to redundancy in the model average (D’Angelo and Fortuin [2021b], Wu et al. [2017], Pang et al. [2019]). This phenomenon of mode collapse is more evident when networks are over-parameterized (D’Angelo and Fortuin [2021b], Kirsch [2024]) and formally been described as the problem of non-identifiability (Roeder et al. [2021]).
@@
The work in this paper addresses the mode collapse issue and proposes a solution to ensure functional diversity using gradient information. It examines how functional diversity in particle optimization methods can be effectively controlled with kernel density estimation to approximate the target density. The paper introduces a simple method for training ensemble members for improved mode exploration and presents empirical results to validate the approach.
@@ 1.1 Contribution
Empirically study the implications of the method and test its efficacy in improving the planning for Model-Based Reinforcement Learning.
2 Deep Ensembles using ERGD
@@ 2.1 Bayesian Neural Networks (BNNs)
In a supervised learning setting, one has where denoting the training inputs and denote the corresponding outputs, with and , respectively.
Let denote a mapping function parameterized by a neural network, and one can define a likelihood function using this .
@@
In Bayesian Neural Networks, the goal is to estimate the posterior distribution that captures all possible explanations for the given training data, defined as follows.
represents the prior distribution over the weight parameters .
Most traditional ways to use neural network methods use a point estimate called a Maximum A Posteriori (MAP) estimate, which can be useful when sampling from a Bayesian Neural Network is challenging or when there is no need to capture model uncertainty.
In this paper, the focus is on capturing a type of uncertainty called epistemic uncertainty. It is the uncertainty associated with estimating the model parameters given the training data. Epistemic uncertainty modeling is well-suited when data is limited or additional data collection is difficult. Although many works use epistemic uncertainty methods when distributional shifts are present, this could be potentially misleading. Ensembles are essential for improving predictive performance, uncertainty quantification, and robustness over a single model. However, their ability to quantify uncertainty and ensure robustness on out-of-distribution data correlates with their in-distribution performance. This falls outside the primary focus for discussion in this paper. But there are non-parametric ways that quantify uncertainty for such distribution shifts (Malinin and Gales [2018]).
@@
However, this integral is intractable for neural networks. In practice, it is approximated using Bayesian Model Averaging (BMA). Sampling-based methods for Bayesian Inference facilitate the acquisition of representative samples from such a posterior distribution to estimate this quantity. This is why prediction accuracy in most approximate methods relies heavily on this Bayesian Model Averaging estimate.
2.2 Particle-optimization Variational Inference
This section describes how Implicit methods estimate the posterior as a stationary distribution by minimizing an energy functional. This stationary distribution results by simulating the system’s dynamics for an infinite amount of time. This can be modeled as a partial differential equation of the form :
where is the approximate posterior at time . is the gradient flow defined as moving in the direction of the steepest descent of an energy or objective function. This gradient flow depends on the choice of metric and energy functional. Samples obtained by simulating such a gradient flow constitute samples from the stationary distribution. In practice, numerical methods can be used for simulation.
Particle-optimization Variational Inference (POVI) methods approximate with a set of particles , i.e.,
simulate the gradient flow with a discretized version of the ordinary differential equation (ODE):
@@ 2.2 Particle-optimization Variational Inference
Implicit Variational methods, thus, allow for the generation of samples without having to do probability evaluation for approximating the true posterior. This contrasts with Explicit variational methods, which approximate the posterior distribution by minimizing the KL divergence between a simpler, tractable distribution and the true posterior. The use of limited tractable distributions can hinder the accuracy of such methods.
Different POVI methods differ in their choices of metric spaces and functionals for gradient flows. A pivotal study that proved significant for subsequent research directions is Stein Variational Gradient Descent (SVGD) (Liu and Liu [2017]). D’Angelo and Fortuin [2021b] shows how SVGD can be interpreted as a gradient flow in the Wasserstein Space under a particular Stein geometry.
Jordan et al. [1998] described the optimization problem of a functional as the evolution in time of the measure under the Wasserstein gradient flow is described by
Here,
@@
For SVGD, the update it given by :
In particular, if is a linear operator defined as and kernel captures the particle interactions (D’Angelo and Fortuin [2021b]), the ODE describing the SVGD update becomes :
Here, is the true posterior and is the target approximation.
The -Wasserstein distance for between and is defines a metric on as
Think of this as an optimal transport problem where one wants to transform elements in the domain of to with a minimum cost.
D’Angelo and Fortuin [2021b] introduces a method that uses Wasserstein gradient flow with kernel density to approximate the target distribution. In their proposed method, the functional is the KL divergence between the target approximation and the true posterior. The update for particle simulation in this work is as follows :
Chen et al. [2018] presents a novel particle-optimization framework grounded in Wasserstein gradient flows to unify two prominent scalable Bayesian sampling methods: stochastic gradient Markov chain Monte Carlo (SG-MCMC) and Stein variational gradient descent (SVGD).
One of the factors impacting the diversity of samples in all the methods mentioned above, obtained at the end of the simulation, is the initial particle distribution. Additionally, as the networks get larger, these samples become less and less sensitive to this initial distribution during training and converge to a single mode. Empirical evidence and theoretical analysis Zhuo et al. [2018] have shown that particles in methods such as SVGD exhibit mode collapse. They converge to a limited number of local modes. It is natural to ask if there is a way to reduce this dependence on the initial distribution and enhance mode coverage for diversity. This is not a novel research question; several prior works have looked into finding ways to answer this question to address mode collapse.
The proposed method, called Entropy-Regularized Gradient Estimators, is a method to enhance mode coverage by mitigating the ensemble’s dependence on the initial distribution while accounting for particle interactions within the ensemble. It uses an empirical measure of particles to approximate the target posterior. This set of particles is taken from an initial distribution and is evolved using the following update :
The full algorithm for Entropy-Regularized Gradient Descent is described in Algorithm 1. A symmetric kernel is used to model the interaction between the particles to train an ensemble. ERGD introduces a regularization constant that controls the effect of this interaction on functional diversity. The key idea behind this is that using kernel density as an approximation to the target density, the constant regularizes cross-entropy between the kernel density and this approximation. Intuitively, the evolution of the particles is influenced more by their interactions with each other.
Particle evolution equation:
where is the particle, is a symmetric positive-definite kernel, and is the particle density.
Total energy functional:
where is a parameter, is the cross-entropy between particle density and kernel density , is the Kullback-Leibler divergence between and target distribution .
@@
The convergence guarantees of the method determine the choice of this regularization constant. As , is a solution of the corresponding Louiville Equation (more details in Appendix A). A simple choice for satisfying this is to use a linear schedule between positive values bigger than 1 to 1. This provides a method to guide particle evolution, placing much greater emphasis on the interactions between the particles in the initial distribution.
Using a symmetric kernel as an approximation to the target probability density changes the gradient flow of KL divergence between two distributions under the metric induced by the Stein Operator to the gradient flow of KL divergence under the Wasserstein Metric (D’Angelo and Fortuin [2021b]). The following update gives the particle evolution method described in D’Angelo and Fortuin [2021b] :
It can be understood as having the ensemble members grow without attraction, leading to a high probability density. In this case, the authors used as 1 / . Such a can be bad in practice (D’Angelo and Fortuin [2021b], D’Angelo and Fortuin [2021b]). Kirsch [2024] developed gradient estimators by inverting Stein’s Identity. The work demonstrates that their method of selecting outperforms the gradient estimator proposed in D’Angelo and Fortuin [2021b], as their estimator accounts for the effects of all interactions on a particle, including those not involving the particle itself. ERGD can direct particle evolution based on an initial distribution, but it’s performance depends on this initial distribution. To find a distribution that optimizes for mode coverage, ERGD is extended to -ERGD as follows :
-ERGD update makes a tunable parameter in the algorithm. Experiments comparing different updates use for -ERGD.
@@ 3.2 Classification Tasks
Different methods are compared for an ensemble size of 10 and 20 for the classification of CIFAR-10 and FMNIST, respectively. Classification on CIFAR-10 is done using ResNet 20 and Fashion-MNIST uses a neural network with 3 hidden layers and 128 hidden units. ERGD and -ERGD are compared with SGD, kde-WGD using an update rule given by [6], sse-WGD as described in Kirsch [2024], and SVGD. ERGD uses a linear annealing schedule. For -ERGD, experiments use a constant value of beta to compare with kde-WGD and sse-WGD.
ERGD performs better than SVGD and is comparable to kde-WGD and sse-WGD for both datasets. Deep Ensembles perform comparably to kde-WGD on CIFAR-10 and have comparative performance to -ERGD on FMINST.
@@
Additionally, the performance of ERGD is compared with SGD as the network size changes. For this, the CIFAR-10 classification task is used, and the size of the ResNet is changed. There are different sizes compared, while the size of the ensemble is the same, i.e., 5. The green plot describes using ERGD with an initial exploratory phase, which does not add any advantage over ERGD for this case. ERGD has higher values of Entropy Ratio and Model disagreement compared to SGD in all cases, while the test accuracy gets close as the size increases.
3.3 Uncertainty-aware Dynamics Models
@@ 4 Related Work
5 Conclusion